OpenStudy (anonymous):

Let f(x) = 9x - 2 and g(x) = -x + 3. Find f(g(x)). -9x - 2 -9x + 5 -9x + 25 -9x + 27

OpenStudy (anonymous):

Any idea how to start?

OpenStudy (anonymous):

Ditribute 9x*-x and 9x*3

OpenStudy (anonymous):

Mhm

OpenStudy (anonymous):

\[f(g(x)) = 9(-x+3)-2\] You just plug g(x) expression where the x is.

OpenStudy (anonymous):

Does that make sense?

OpenStudy (anonymous):

Now just distribute.

OpenStudy (anonymous):

Yes, I got the answer thankyou batman!

OpenStudy (anonymous):

Np ^.^

OpenStudy (ciarán95):

Let's take the function f(x) = 9x -2. If we wanted to find f(2), we would sub in '2' to wherever we found x in the function (i.e. we would get 9(2) - 2 = 18 - 2 = 16). So, if we were looking for f(g(x)), we would sub in g(x), or -x + 3, into everywhere we found x. So, we would get f(g(x)) = 9(-x + 3) -2, which you can simplify to give you the answer.